OpenStudy (freemap):

Joel uses recycled materials to build and sell art sculptures. For s small sculptures, his revenue is modeled by r(s) = 45s and his cost is modeled by c(s) = 10 + 3.5s. Which function is Joel’s profit function? p(s) = 10 – 41.5s p(s) = 10 + 48.5s p(s) = 41.5s – 10 p(s) = 48.5s – 10

OpenStudy (freemap):

@ganeshie8

ganeshie8 (ganeshie8):

quick question, do you know what the term `revenue` means ?

OpenStudy (freemap):

There income right?

ganeshie8 (ganeshie8):

Kindof, revenue is the monety that comes from Sales

OpenStudy (freemap):

Thats what I meant

ganeshie8 (ganeshie8):

Right, so how to find `profit` function when they give you `revenue` and `cost` functions ?

OpenStudy (freemap):

I'm kind of confused because i was thinking just add. But are they talking about the cost that he had to take out, materials to buy. Then it says he uses recycled material. Therefore he doesn't spend any money of buying materials. So would you add both rev. and cost? I don't think you do, but just asking

ganeshie8 (ganeshie8):

Very good, you're paying attention to details, which is nice :) I think we should subtract the cost because he may be paying building rent / electricity bills right ?

OpenStudy (freemap):

Thank you, oh okay so cost=10+3.5 =13.5 and of rev= 45s. Subtract 45-13.5=31.5 which is not a choice,. 45-10=35 not an option. 4-10+3.5=38.5 also not an option. Wait 45-3.5 is 41.5 10- 41.5?

ganeshie8 (ganeshie8):

Remember this phrase when you're in algebra : we cannot add coconuts and oranges

ganeshie8 (ganeshie8):

when we count number of coconuts and oranges in a bag, we add coconuts to coconuts we add oranges to oranges

ganeshie8 (ganeshie8):

In algebra also, we only add `like terms` together

ganeshie8 (ganeshie8):

r(s) = 45s c(s) = 10 + 3.5s p(s) = r(s) - c(s) = 45s - (10 + 3.5s) = 45s - 10 - 3.5s

ganeshie8 (ganeshie8):

you need to treat the terms containint "s" as a different species

ganeshie8 (ganeshie8):

you can't subtract 45s and 10 ok

ganeshie8 (ganeshie8):

*containing

OpenStudy (freemap):

45 and 3.5 s are the like terms therefore I add 45+3.5=48.5s Then the 10, I just say 10-48.5s that would be dividing. Because of that s

ganeshie8 (ganeshie8):

Still wrong, we need to subtract cost from the revenue

OpenStudy (freemap):

the cost is 10 ok 10 - 48.5 is 38.5

ganeshie8 (ganeshie8):

Nope, start over

ganeshie8 (ganeshie8):

r(s) = 45s c(s) = 10 + 3.5s First, can you write a statement for the profit, p(s) ?

OpenStudy (freemap):

p/s ratio?

ganeshie8 (ganeshie8):

no, you have said earlier that profit equals revenue minus cost right ?

OpenStudy (freemap):

yes, but what i keep subtracting isn't right. So I don't know what to do

ganeshie8 (ganeshie8):

lets go step by step, profit = revenue - cost = 45s - (10 + 3.5s) fine till here ?

ganeshie8 (ganeshie8):

i have just substitute the expressions for revenue and cost

OpenStudy (freemap):

45s - (10+3.5s ) Like you said I cant ad apples and oranges I can't add 10+3.5 because of the s So if I do okay 45s - (13.5)=31.5

ganeshie8 (ganeshie8):

you can't add 10 and 3.5s but you can surely combine 45s and 3.5s

ganeshie8 (ganeshie8):

let me just show you next step

ganeshie8 (ganeshie8):

maybe let me just give you the complete solution and see if you follow it

OpenStudy (freemap):

I did add 45s and 3.5 =48.5 I thought you said that was wrong

ganeshie8 (ganeshie8):

r(s) = 45s c(s) = 10 + 3.5s p(s) = r(s) - c(s) = 45s - (10 + 3.5s) = 45s - 10 - 3.5s = (45s - 3.5s) - 10 = 41.5s - 10

OpenStudy (freemap):

I had 41.5 -10 Up top

OpenStudy (freemap):

I get why you said it was wrong I had 10-41.5 not 41.5-10

OpenStudy (freemap):

C Right?

OpenStudy (freemap):

Thank you

OpenStudy (rational):

C is correct! nice job staying till the end :)